# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v1

@@ B.3. Contractivity
    Commutation with constants: If $F\in CDF(\overline{\mathbb{R}})$ satisfies $c_{-}\leq F\leq c_{+}$ for some constants $c_{-},c_{+}\in[0,1]$, then $T(F+c)=TF$ for any $c\in[-c_{-},1-c_{+}]$.
@@
    Stationarity at infinity: For any $F\in CDF(\overline{\mathbb{R}})$, $TF(+\infty)=F(+\infty)$ and $TF(-\infty)=F(-\infty)$.
Then $T$ is contractive, that is,
$\|TF-TG\|_{\sup}\leq\|F-G\|_{\sup}\quad\text{for any}\,\,F,G\in CDF(\overline{\mathbb{R}}).$
The proof is based on the following lemma.
@@
Let $T$ be an operator on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(iii) of Proposition 12. If $F_{1},F_{2}\in CDF(\overline{\mathbb{R}})$ are chosen so that
$F_{2}\leq TF_{1}\quad\text{pointwise in}\,\,\mathbb{R}$,
then, for any $\delta\in[0,1]$ and $c\in[0,1-\delta]$, the functions $G_{1},G_{2}$ defined by
$G_{i}=\max\{F_{i}-\delta,c\}$
satisfy $G_{2}\leq TG_{1}$ pointwise in $\mathbb{R}$.
@@
Notice that $G_{1}+\delta=\max\{F_{1},c+\delta\}$, which is a function in $CDF(\overline{\mathbb{R}})$ that is pointwise larger than $F_{1}$. Therefore, by monotonicity,
$T(G_{1}+\delta)\geq TF_{1}\geq F_{2}.$
On the other hand, since property (iii) implies $T(G_{1}+\delta)(-\infty)=G_{1}(-\infty)+\delta\geq c+\delta$, one deduces that $T(G_{1}+\delta)\geq c+\delta$ pointwise. Therefore, $T(G_{1}+\delta)\geq\max\{F_{2},c+\delta\}=G_{2}+\delta$ follows. By property (ii), this implies $TG_{1}\geq G_{2}$. ∎
Here is the proof of contractivity:
Proof of Proposition 12.
Let $F,G\in\text{CDF}(\overline{\mathbb{R}})$. Note that
$G-\|F-G\|_{\text{sup}}\leq F\leq G+\|F-G\|_{\text{sup}}$
and the goal is to show that the same inequality holds if $(F,G)$ is replaced by $(TF,TG)$. By symmetry, it suffices to prove that $TF\leq TG+\|F-G\|_{\text{sup}}$.
Since the desired inequality is trivial if $\|F-G\|_{\text{sup}}\geq 1$, assume that $\|F-G\|_{\text{sup}}<1$ in what follows.
Let $G_{1}=\max\{F-\|F-G\|_{\text{sup}},0\}$ and $G_{2}=\max\{TF-\|F-G\|_{\text{sup}},0\}$. By Lemma 4,
$G_{2}\leq TG_{1}.$
Yet $G_{1}\leq G$ holds so $G_{2}\leq TG_{1}\leq TG$ by (i). This implies $TF-\|F-G\|_{\text{sup}}\leq TG$.∎
@@ Appendix C Effective Coefficients in the Series-Parallel Graph
These functions have the power series representation $Li_{k}(t)=\sum_{n=1}^{\infty}n^{-k}t^{n}$, hence, in particular, $Li_{k}(1)=\zeta(k)$. The computations that follow seem to be related to Euler’s formula for $Li_{2}(\frac{1}{2})$ and Landen’s formula for $Li_{3}(\frac{1}{2})$ (see [19, Section 1.1]).

### v2

@@ B.3. Lipschitz Estimate
    Discrete-Time Reaction-Diffusion Equation: There is a function $\mathcal{L}$ defined on $CDF(\overline{\mathbb{R}})$ and a function $Q$ defined on $[0,1]$ such that
    $TF-F=\mathcal{L}F-Q(F)\quad\text{for each}\,\,F\in CDF(\overline{\mathbb{R}}).$
@@
    Commutation with Constants: Given $F,G\in CDF(\overline{\mathbb{R}})$, if the difference FGF-G is a constant function, then $\mathcal{L}F=\mathcal{L}G$.
If there is a constant $L>0$ such that $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|$ for each $q,q^{\prime}\in[0,1]$, then
$\|T^{n}F-T^{n}G\|_{\sup}\leq(1+L)^{n}\|F-G\|_{\sup}\quad\text{for any}\,\,F,G\in CDF(\overline{\mathbb{R}}),\,\,n\in\mathbb{N}.$
The proposition follows immediately from the following lemma, which was needed in the proof of Theorem 3 above:
@@
Suppose that $T$ is a monotone operator on $CDF(\overline{\mathbb{R}})$ satisfying the assumptions of Proposition 17. If $\{F_{n}\}$ is any sequence with $F_{n}\leq TF_{n-1}$ for each $n$, then, for any $\delta>0$, the sequence $\{\tilde{F}_{n}\}$ given by $\tilde{F}_{n}=\max\{F_{n}-\delta(1+L)^{n},0\}$ satisfies
(92) $\tilde{F}_{n}\leq T\tilde{F}_{n-1}\quad\text{for each}\,\,n.$
In particular, if $\{F_{n}\}$ and $\{G_{n}\}$ are any two sequences in $CDF(\overline{\mathbb{R}})$ such that
$F_{n}\leq TF_{n-1}\quad\text{and}\quad G_{n}\geq TG_{n-1}\quad\text{ for each}\,\,n\in\mathbb{N},$
then
$F_{n}\leq G_{n}+(1+L)^{n}\sup\left\{(F_{0}(x)-G_{0}(x))_{+}\,\mid\,x\in\mathbb{R}\right\}\quad\text{for each}\,\,n.$
@@
First, consider the sequence $\{\tilde{F}_{n}\}$. Let $\{c_{n}\}_{n\in\mathbb{N}}$ be the sequence of nonnegative numbers determined by the recursion
$c_{n}=(1+L)c_{n-1},\quad c_{0}=\delta,$
so that $\tilde{F}_{n}=\max\{F_{n}-c_{n},0\}$.
Fix $n\in\mathbb{N}$. If $c_{n-1}\geq 1$, then (92) holds trivially as $\tilde{F}_{n}=\tilde{F}_{n-1}=0$ in this case. Thus, assume $c_{n-1}\in(0,1)$. By monotonicity and the assumption on $\mathcal{L}$,
$TF_{n-1}-c_{n}$ $\leq T(\max\{F_{n-1},c_{n-1}\})-c_{n}$
$=\max\{F_{n-1},c_{n-1}\}+\mathcal{L}(\max\{F_{n-1}-c_{n-1},0\})-Q(\max\{F_{n-1}-c_{n-1},0\})$
$\qquad-Q(\max\{F_{n-1},c_{n-1}\})+Q(\max\{F_{n-1}-c_{n-1},0\})-c_{n}$
$\leq T(\max\{F_{n-1}-c_{n-1},0\})+c_{n-1}-c_{n}+Lc_{n-1}.$
In view of the formula for $\{c_{n}\}$ and the property of $\{F_{n}\}$, this implies
$F_{n}-c_{n}\leq T(\max\{F_{n-1}-c_{n-1},0\})=T\tilde{F}_{n-1}.$
In view of the fact that $TG\geq 0$ for any $G\in CDF(\overline{\mathbb{R}})$, this last bound implies (92).
Finally, suppose that, in addition, $G_{n}\geq TG_{n-1}$ for any $n$, and let $\delta=\sup\{(F_{0}(x)-G_{0}(x))_{+}\,\mid\,x\in\mathbb{R}\}$. Notice that $\tilde{F}_{0}=\max\{F_{0}-c_{0},0\}\leq G_{0}$ by definition. Thus, by the monotonicity of $T$,
$\max\{F_{n}-c_{n},0\}=\tilde{F}_{n}\leq G_{n}\quad\text{for each}\,\,n.$
In particular, $F_{n}\leq G_{n}+c_{n}$ for any $n$, as desired. ∎
@@ Appendix C Effective Coefficients in the Series-Parallel Graph
These functions have the power series representation $Li_{k}(t)=\sum_{n=1}^{\infty}n^{-k}t^{n}$, hence, in particular, $Li_{k}(1)=\zeta(k)$. The computations that follow seem to be related to Euler’s formula for $Li_{2}(\frac{1}{2})$ and Landen’s formula for $Li_{3}(\frac{1}{2})$ (see [22, Section 1.1]).